OpenStudy (anonymous):

–3v – 2v > 35 solve for v

jimthompson5910 (jim_thompson5910):

what does –3v – 2v simplify to?

OpenStudy (anonymous):

1v..?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

-3 - 2 = _____

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

think of it like: you're 3 dollars in debt and you spend 2 more dollars

jimthompson5910 (jim_thompson5910):

yep, -3 - 2 = -5

jimthompson5910 (jim_thompson5910):

so -3v - 2v = -5v

jimthompson5910 (jim_thompson5910):

which means –3v – 2v > 35 turns into –5v > 35

OpenStudy (anonymous):

Divide both sides..?

jimthompson5910 (jim_thompson5910):

the goal is to isolate v, so how do we do that? how do we move that -5 over?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

by what

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

correct, so that will flip the sign (because you're dividing both sides by a negative number)

jimthompson5910 (jim_thompson5910):

–5v > 35 –5v/(-5) < 35/(-5) v < -7

OpenStudy (anonymous):

thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome